OpenStudy (shae.mccoy):

A circle with area 81pi has a sector with a central angle of 310 degrees. What is the area of the sector?

OpenStudy (shae.mccoy):

@gorv

OpenStudy (gorv):

circle measure 360degree agree??

OpenStudy (gorv):

2 *pi measure 81 pi 1 will measure =81pi/2Pi

OpenStudy (shae.mccoy):

hole on

OpenStudy (shae.mccoy):

hold on

OpenStudy (gorv):

310 degree we need to convert it into radian

OpenStudy (shae.mccoy):

ok

OpenStudy (shae.mccoy):

@gorv

OpenStudy (shae.mccoy):

where'd you go?

OpenStudy (gorv):

u know isosceles triangle ??

OpenStudy (gorv):

lol i m asking u ??

OpenStudy (shae.mccoy):

Oh.. ok yes

OpenStudy (shae.mccoy):

@Ahaanomegas

OpenStudy (mathstudent55):

\(\large A_{sector} = \dfrac{\theta}{360}\pi r^2\) Where \(\theta\) is the measure of the central angle of the sector in degrees.

OpenStudy (mathstudent55):

In the formula above, \(\pi r^2\) is the area of the circle. In this problem, you know the area of the circle is \(81 \pi\), so just replace \(\pi r^2\) with \(81\pi\). Then replace \(\theta\) with 310.

OpenStudy (mathstudent55):